# Pretrained Reversible Generation as Unsupervised Visual Representation Learning

### v1

@@ 4.3 Advantages
Compared to previous approaches, our method offers several key advantages:
@@
Our approach is agnostic to the choice of neural network architecture for the generative model — it can be a U-Net, Transformer, or any other model. The latent variable $Z$ remains stationary with respect to the data $X$ when using an ODE solver due to the static nature of continuous-time stochastic flow models, as observed by song2020score. This allows different architectures to encode the data while yielding the same latent variable $Z$. In contrast, methods like DDAE [xiang2023denoising] are model-dependent, with the latent variable $Z$ tied to specific network architectures and derived from intermediate activations, making them less flexible and non-stationary.
@@
Our method leverages the infinite-layer structure of continuous-time stochastic flow models [Chen2018], providing high expressiveness with relatively small parameter sizes. This structure underpins the success of modern generative models like Stable Diffusion 3 [rombach2022high] and DALL·E 3 [betker2023improving]. By using a reversible process, our method allows for the simultaneous training of discriminative and generative models, offering state-of-the-art performance with this enhanced capacity.
@@
Features extracted by reversing the generative process exhibit stability along the reverse trajectory, enabling the use of features from any point $t\in[0,1]$ for downstream tasks, not just the latent variable $x_{1}$. These features are robust to different discretization schemes (e.g., Euler, Runge-Kutta) and various time steps (ranging from $10^{3}$ to $10^{1}$). Additionally, our method generalizes well across datasets, meaning that a powerful pretrained model can be fine-tuned further for improved performance on new tasks.
In summary, our approach bridges the gap between generative and discriminative models, offering a simple yet effective way to leverage pretrained generative models for unsupervised representation learning, with the ability to fine-tune for downstream supervised tasks.
@@ 5 Experiments
We conducted experiments to evaluate our method across three key aspects. First, we analyzed the effectiveness of pretraining and finetuning strategies to identify critical components of our approach. Second, we benchmarked our method against state-of-the-art generative and classical classification methods on CIFAR-10 [Krizhevsky2009LearningML], Tiny-ImageNet [Le2015TinyIV], and ImageNet [Russakovsky2014ImageNetLS]. Finally, we performed ablation studies to assess the impact of different design choices—such as architecture and optimization techniques—on downstream classification performance.
@@ 5.1 Settings
For we use three diffusion and flow models, GVP, ICFM, and OTCFM. For pretraining, we followed the protocol from [nichol2021improved], using the Adam optimizer [kingma2014adam] with a fixed learning rate of $1\times 10^{-4}$ for 1,200 epochs on the respective training sets. For downstream image classification, we adopted the configuration from [liu2021swin], using AdamW for 200 epochs with a cosine decay learning rate schedule and a 5-epoch linear warm-up. The training used a batch size of 128 and an initial learning rate of 0.001. To ensure a controlled setting, we excluded advanced regularization techniques like mixup [zhang2017mixup] and cutmix [yun2019cutmix], relying on simple data augmentation. Further details, including hyperparameters and computation efficiency, are provided in Secs. B.1 and B.2 .
@@ 5.2 Verification Analysis
We aim to address the following four questions.
5.2.1 Better Pretraining Lead to Better Finetuning?
To explore whether better pretraining improves finetuning, we compared the classification accuracies of models finetuned from different pretraining epochs on CIFAR-10 (Fig. 4). We also computed the mutual information during pretraining, which increases monotonically, as shown in Fig. 4, and analyzed the relationship between accuracy and mutual information during finetuning in Fig. 4.
Models without pretraining achieved approximately $73.5\%$ accuracy. In contrast, the more pretraining a model received, the higher its mutual information and classification accuracy, indicating that stronger generative capability leads to better finetuning performance. During finetuning, we observed a decline in $\mathcal{I}(x_{0},x_{1})+\mathcal{H}(x_{1})-\mathcal{H}(x_{0})$ for both the training and validation sets, suggesting that the model filters out unnecessary features to improve downstream classification, which can also be observed in Fig. 5 as some visual representations are kept while others are discarded. In summary, sufficient pretraining is crucial for optimal performance in inverse generative classification.
5.2.2 What is a Reasonable Finetuning Strategy?
Figure 5: Reverse generation from $x_{0}$ to $x_{1}$ using fully pretrained and fine-tuned models. See Sec. B.4 for details.
@@ 5.2.2 What is a Reasonable Finetuning Strategy?
We first examine whether a frozen flow model provides meaningful features. For this, we use features $x_{t}$ along the sampling trajectory from $x_{0}$ to $x_{1}$ under two conditions: (1) freezing the generative model’s parameters (light lines) and (2) end-to-end training of both the generative model and the classification head (dark lines). As shown in Fig. 6, the performance gap is substantial—47.10% on CIFAR-10 and 58.04% on Tiny-ImageNet—demonstrating the critical importance of updating the generative model’s parameters during finetuning.
Additionally, on CIFAR-10, initiating finetuning from later stages of the sampling trajectory ($x_{1/4}$ to $x_{1}$) results in progressively better performance. On TinyImageNet, optimal results are achieved when finetuning starts from $x_{1/2}$ to $x_{1}$. This difference likely stems from the complexity of the datasets: more complex datasets require stronger feature extraction, necessitating longer trajectory lengths for effective finetuning. Further experiments exploring optimal finetuning strategies are detailed in the ablation studies.
5.2.3 Does the Model Serve as a Continuous Feature Extractor?
Residual networks, such as ResNet-50 [he2016deep], achieve feature extraction by constructing discrete-time methods through a series of composite transformations: $\mathbf{h}_{t+1}=\mathbf{h}_{t}+f(\mathbf{h}_{t},\theta_{t})$. In contrast, we construct the feature extractor using the ODE specified by the neural network, namely $\frac{d\mathbf{h}(t)}{dt}=f(\mathbf{h}(t),t,\theta)$. Tabs. 1 and 2 demonstrate that although we trained using a discrete sample length of $t_{\text{span}}=20$, we can evaluate the model using any reasonable inference steps. Furthermore, feature extraction is not limited to a specific training point (such as the midpoint); extracting features within its vicinity (within 20%) still achieves high performance on downstream tasks. These findings validate our method can serve as a continuous feature extractor.
@@ 5.2.3 Does the Model Serve as a Continuous Feature Extractor?
WideResNet-$28$-$10$ [zagoruyko2016wide] $36$ $96.3$
ResNeXt-$29$-$16*64$d [xie2017aggregated] $68$ $96.4$
@@
GLOW [fetaya2019understanding] N/A $84.0$
Energy model [duvenaud2020your] N/A $92.9$
SBGC [zimmermann2021score] N/A $95.0$
HybViT [yang2022your] $43$ $96.0$
DDAE [xiang2023denoising] $36$ $97.2$
@@
HybViT [yang2022your] $43$ $56.7$
DDAE [xiang2023denoising] $40$ $69.4$
@@
Supplementary Material
Appendix A Diffusion process definition
Full definition of diffusion process that studied in this paper is a Generalized VP-SDE (GVP) using triangular function as coefficients [Albergo2023]:
$\alpha_{t}=\cos\left(\frac{1}{2}\pi t\right),$ (A1)
$\sigma_{t}=\sin\left(\frac{1}{2}\pi t\right).$

### v2

@@ 3.3 Advantages
Our approach is agnostic to the choice of network architecture for the generative model: it can be a U-Net, Transformer, or any other model. The latent variable $Z$ remains stationary with respect to the data $X$ when using an ODE solver due to the static nature of continuous-time stochastic flow models [56]. This allows different architectures to encode the data while yielding the same latent variable $Z$. In contrast, methods like DDAE [65] are model-dependent, with the latent variable $Z$ tied to specific network architectures and derived from intermediate activations, making them less flexible and non-stationary.
@@
PRG leverages the infinite-layer structure of continuous-time stochastic flow models [11], providing high expressiveness with relatively small parameter sizes. This structure underpins the success of modern generative models like SD3 [52] and DALL·E 3 [7]. By using a reversible process, our method allows for the simultaneous training of discriminative and generative models, offering excellent performance with enhanced capacity.
@@
Features extracted by reversing the generative process remain stable along the reverse trajectory, allowing feature use from any point $t\in[0,1]$ for downstream tasks, not just $x_{1}$. These features are robust to various discretization schemes (e.g., Euler, Runge-Kutta) and time steps ($10^{3}$ to $10^{1}$). Our method generalizes well across datasets, enabling fine-tuning of pretrained models for new tasks. Notably, it applies to all diffusion and flow models and has been successfully transferred to community-developed text-to-image models in Sec. 4.4.3.
In summary, our approach bridges the gap between generative and discriminative models, offering a simple yet effective way to leverage pretrained generative models for unsupervised representation learning, with the ability to fine-tune for downstream supervised tasks. A more detailed analysis of why PRG is effective can be found in Appendix A.
@@ 4.1 Settings
For generative model selection, we adopt three diffusion and flow-based models: GVP, ICFM, and OTCFM. For pre-training, we followed the protocol from [48], using the Adam optimizer [30] with a fixed learning rate of $1\times 10^{-4}$ for 1,200 epochs on the respective training sets. For downstream image classification, we adopted the configuration from [39], using AdamW for 200 epochs with a cosine decay learning rate schedule and a 5-epoch linear warm-up. The training used a batch size of 128 and an initial learning rate of 0.001. We followed DDAE [65]’s setting (64 resolution, simple data augmentation, no mixup [71], no cutmix [69]) for fair comparison, yet our method achieves SOTA diffusion classifier performance. Further details, including hyperparameters and efficiency statistics for training and inference, are provided in Secs. B.1, B.2 and B.3 .
4.2 baselines
We select state-of-the-art (SOTA) representation models as baselines to evaluate PRG on CIFAR-10 [32], Tiny-ImageNet [33], and ImageNet [53]. Baselines include classical discriminative models like WideResNet-$28$-$10$[70] and ResNeXt-$29$-$16\times 64$d[66], as well as diffusion-based generative models such as GLOW [19], Energy Model [18], SBGC [73], HybViT [67], and DDAE [65]. For out-of-distribution (OOD) evaluation, we compare PRG with PGD [43], PLAT [64], AugMix [25], AutoAug [15], SBGC, and PDE+ [68] on CIFAR-10-C and Tiny-ImageNet-C. All results are from their original papers.
@@ 4.3 Verification Analysis
4.3.1 Better Pre-training Lead to Better Fine-tuning?
To explore whether better pre-training improves fine-tuning, we compared the classification accuracies of models finetuned from different pre-training epochs on CIFAR-10 (Fig. 4). We also computed the mutual information during pre-training, which increases monotonically, as shown in Fig. 4. During pre-training, we have
$\mathbb{E}_{p(x_{1},x_{0})}\bigl{[}\log p(x_{0}\mid x_{1})\bigr{]}=\mathcal{I}(X_{0},X_{1})\;-\;\mathcal{H}(X_{0}),$
where $\mathcal{H}(X_{0})$ is the (constant) entropy of the data distribution and $\mathcal{H}(X_{1})$ is the (constant) entropy of the Gaussian distribution. Hence, $\mathbb{E}_{p(x_{1},x_{0})}[\log p(x_{0}\mid x_{1})]$ faithfully reflects the mutual information $\mathcal{I}(X_{0},X_{1})$ during pre-training. In contrast, during fine-tuning, $\mathbb{E}_{p(x_{1},x_{0})}[\log p(x_{0}\mid x_{1})]$ no longer represents $\mathcal{I}(X_{0},X_{1})$, because $\mathcal{H}(X_{1})$ is no longer the entropy of a Gaussian distribution. Nonetheless, the quantity $\mathcal{I}(X_{0},X_{1})+\mathcal{H}(X_{1})\;-\;\mathcal{H}(X_{0})$ still captures the change in $\mathcal{I}(X_{0},X_{1})$ to a large extent. We compute log-probability density $\log p(x_{0}|x_{1})$ using the adjoint method or Neural ODEs [11] for computational feasibility:
$\begin{bmatrix}x_{0}-x_{1}\\ \log{p(x_{0})}-\log{p(x_{1})}\end{bmatrix}=\int_{t=1}^{t=0}\begin{bmatrix}v_{\theta}(x_{t},t)\\ -\text{Tr}\left(\frac{\partial v_{\theta}}{\partial x_{t}}\right)\end{bmatrix}\mathrm{d}t$ (10)
For efficient computation of $\log p(x_{0}|x_{1})$, we employ the Hutchinson trace estimator [28, 22]:
$\text{Tr}\left(\frac{\partial v_{\theta}}{\partial x_{t}}\right)=E_{p(\epsilon)}[\epsilon^{T}\frac{\partial v_{\theta}}{\partial x_{t}}\epsilon]$ (11)
where $\epsilon$ is a standard Gaussian random vector.
Moreover, we analyzed the relationship between accuracy and mutual information during fine-tuning, as shown in Fig. 4. Models without pre-training achieved approximately $73.5\%$ accuracy. In contrast, models with more pre-training exhibit higher mutual information and classification accuracy, suggesting that stronger generative capability enhances fine-tuning performance. During fine-tuning, we observed a decline in $\mathcal{I}(x_{0},x_{1})+\mathcal{H}(x_{1})-\mathcal{H}(x_{0})$ for both the training and validation sets, suggesting that the model filters out unnecessary features to improve downstream classification, which can also be observed in Fig. 5 as some visual representations are kept while others are discarded. In summary, sufficient pre-training is crucial for optimal performance in inverse generative classification.
4.3.2 What is a Reasonable Fine-tuning Strategy?
Figure 5: Reverse generation from $x_{0}$ to $x_{1}$ with full pre-training and fine-tuning. See Appendix D for more details.
@@ 4.3.2 What is a Reasonable Fine-tuning Strategy?
We first examine whether a frozen flow model provides meaningful features. For this, we use features $x_{t}$ along the sampling trajectory from $x_{0}$ to $x_{1}$ under two conditions: (1) freezing the generative model’s parameters (light lines) and (2) end-to-end training of both the generative model and the classification head (dark lines). As shown in Fig. 6, the performance gap is substantial—47.10% on CIFAR-10 and 58.04% on Tiny-ImageNet—demonstrating the critical importance of updating the generative model’s parameters during fine-tuning.
Additionally, on CIFAR-10, initiating fine-tuning from later stages of the sampling trajectory ($x_{1/4}$ to $x_{1}$) results in progressively better performance. On TinyImageNet, optimal results are achieved when fine-tuning starts from $x_{1/2}$ to $x_{1}$. This difference likely stems from the complexity of the datasets: more complex datasets require stronger feature extraction, necessitating longer trajectory lengths for effective fine-tuning. Further experiments exploring optimal fine-tuning strategies are detailed in the ablation studies.
4.3.3 Is the Model a Continuous Feature Extractor?
Residual networks, such as ResNet-50 [23], achieve feature extraction by constructing discrete-time methods through a series of composite transformations: $\mathbf{h}_{t+1}=\mathbf{h}_{t}+f(\mathbf{h}_{t},\theta_{t})$. In contrast, we construct the feature extractor using the ODE specified by the neural network, namely $\frac{d\mathbf{h}(t)}{dt}=f(\mathbf{h}(t),t,\theta)$. Tabs. 1 and 2 demonstrate that although we trained using a discrete sample length of $t_{\text{span}}=20$, we can evaluate the model using any reasonable inference steps. Furthermore, feature extraction is not limited to a specific training point (such as the midpoint); extracting features within its vicinity (within 20%) still achieves high performance on downstream tasks. These findings validate our method can serve as a continuous feature extractor.
@@ 4.3.3 Is the Model a Continuous Feature Extractor?
WideResNet-$28$-$10$ [70] $36$ $96.3$
ResNeXt-$29$-$16*64$d [66] $68$ $96.4$
@@
GLOW [19] N/A $84.0$
Energy model [18] N/A $92.9$
SBGC [73] N/A $95.0$
HybViT [67] $43$ $96.0$
DDAE [65] $36$ $97.2$
@@
HybViT [67] $43$ $56.7$
DDAE [65] $40$ $69.4$
@@
Method Param. (M) Acc. (%)
Discriminative methods
ViT-L/$16$ ($384^{2}$) [17] $307$ $76.5$
ResNet-$152$ ($224^{2}$) [23] $60$ $77.8$
Swin-B ($224^{2}$) [39] $88$ $83.5$
Generative methods
HybViT ($32^{2}$) [67] $43$ $53.5$
DMSZC-DiTXL2 ($256^{2}$) [34] $338$ $77.5$
iGPT-L ($48^{2}$) [10] $1362$ $72.6$
[…]
PRG-GVP-onlyPretrain ($64^{2}$) $122$ $20.18$
PRG-GVP-XL ($64^{2}$) $122$ $77.84$
PRG-ICFM-XL ($64^{2}$) $122$ $78.12$
PRG-OTCFM-XL ($64^{2}$) $122$ $78.13$
Table 5: Comparison on the ImageNet dataset with various algorithms. All results are reported from their original papers, the values in parentheses indicate the input image size.
4.4 Main Results
4.4.1 Performance on Image Classification
To comprehensively evaluate our method, we conducted experiments on three image classification datasets, as detailed in Secs. 4.3.3, 4.3.3 and 4.3.3. Our methods achieved accuracies of 97.59%, 71.12%, and 78.1% on CIFAR-10, Tiny-ImageNet, and ImageNet, respectively, surpassing all existing generative approaches, as well as supervised methods like WideResNet-28[70]. However, these results still trail behind recent transformer-based supervised learning architectures like SwinTransformer [39]. We hypothesize that integrating transformer architectures and leveraging VAEs to handle high input resolutions could further enhance the performance. This avenue will be explored in future work.
4.4.2 Out-of-Distribution Robustness
To address the degradation in out-of-distribution (OOD) due to common image corruptions or adversarial perturbations [73], data augmentations and adversarial training are typically employed. However, recent studies [73, 8, 34] have indicated that generative-based classifiers, without requiring additional data, often exhibit superior OOD robustness. As illustrated in Sec. 4.4.2 and Sec. C.3, our method showcases remarkable robustness on these two datasets with only simple data augmentation.
Model CIFAR-10 CIFAR-10-C
Clean Corr Severity All w/ Noise
Adversarial Training
PGD *[43] $93.91$ $83.08$ $\downarrow 10.83$ $82.10$ $\downarrow 11.81$
PLAT * [64] $94.73$ $88.28$ $\downarrow 6.45$ 88.5688.5688.5688.566.17absent6.17$\downarrow$ 6.17↓ 6.17
Noise Injection
RSE *[37] $95.59$ $77.86$ $\downarrow 17.73$ N/A
ENResNet *[64] $83.33$ $74.34$ $\downarrow 8.99$ N/A
Data Augment
AugMix § [25] $95.83$ $89.09$ $\downarrow 6.74$ $88.71$ $\downarrow 7.12$
AutoAug *[15] $95.61$ $85.37$ $\downarrow 10.24$ 76.4776.4776.4776.4719.14absent19.14$\downarrow$ 19.14↓ 19.14
Generative Methods
SBGC [73] $95.04$ $76.24$ $\downarrow 18.80$ 75.3875.3875.3875.3819.66absent19.66$\downarrow$ 19.66↓ 19.66
PDE+ [68] $95.59$ $89.11$ $\downarrow 6.48$ 85.5985.5985.5985.5910.00absent10.00$\downarrow$ 10.00↓ 10.00
[…]
PRG-GVP-S $97.35$ $91.21$ $\downarrow 6.14$ $92.17$ $\downarrow 5.18$
PRG-ICFM-S $97.59$ $92.13$ $\downarrow 5.46$ $93.07$ $\downarrow 4.52$
PRG-OTCFM-S $97.65$ $92.26$ $\downarrow 5.39$ $93.10$ $\downarrow 4.55$
Table 6: (OOD: Image Corruptions) Comparison on CIFAR-10-C, including Noise, Blur, Weather, and Digital corruptions. Results are referenced from original papers or [68, 73]. * denotes ResNet-18 as the base model, while § indicates ResNeXt-29.
Method Param. (M) Cifar Acc. Tiny Acc.
Classic self-supervised learning methods
MAE (ViT-B/16) [40] $86$ $96.5$ $76.5$
SimCLR Res-50-4x [12] $375$ $98.6$ N/A
Generative methods
DDAE-DiT-XL2 [65] $338$ $98.4$ $77.8$
iGPT-L [10] $1362$ $99.0$ N/A
[…]
PRG-SiT-XL2 $338$ $98.72$ $78.33$
Table 7: Comparison of models pretrained on ImageNet-1k and transferred (fine-tuning) to CIFAR-10 and Tiny-ImageNet.
4.4.3 Transferring Features
A main goal of the pre-training/fine-tuning paradigm is to learn transferrable features. We hypothesize that advancements like powerful pretrained models in the generative model community can contribute to our method. To validate this, we conducted experiments by transferring the pretrained SiT-XL [42] model to two other datasets. Since SiT provides only class-conditional checkpoints, we adopted an unconditional approach by setting the label parameter to null. Specifically, we fine-tuned the model for 28 epochs on the CIFAR-10 dataset and for 45 epochs on the Tiny-ImageNet dataset. As demonstrated in Sec. 4.4.2, our method shows superior transfer learning performance. The results suggest that our algorithm benefits from larger datasets and the latent space of generative models.
4.5 Ablation Studies
In this section, we conduct three ablation studies to further evaluate the applicability of our method. Additional experiments, including analyses of loss type, ODE solver choice, dual-task supervision, and $t_{\text{span}}$, are provided in Appendix C.
4.5.1 Comparison with Generation without pre-training
We compare PRG with classifiers trained without generative pre-training. To ensure fairness, we train the latter model for 600 epochs until its performance no longer improves. As shown in Tab. 8, PRG without pre-training, relying solely on a single supervision signal, may discard useful information, leading to suboptimal performance. This highlights that the intermediate latent variables of a pretrained flow model provide a strong initialization, requiring only slight fine-tuning for optimal feature extraction.
CIFAR-10 Tiny-ImageNet ImageNet
PRG w/o pre-training $0.70$ $0.35$ $0.42$
PRG $0.97$ $0.71$ $0.78$
Table 8: Effect of pre-training: Performance comparison of PRG with and without pre-training on the CIFAR-10 and Tiny-ImageNet.
4.5.2 Effect of $\beta$ on Accuracy
During pre-training, we introduced $\beta$ in Eq. 3 to enhance the mutual information lower bound while balancing the downstream task objective. Table 9 shows PRG scores across datasets for different $\beta$ values.
$\beta$ 1 10 100
Acc. (CIFAR/Tiny) $0.973$/$0.712$ $0.974$/$0.711$ $0.965$/$0.673$
Table 9: ($\beta$ values) Accuracy scores of PRG with varying $\beta$ values on the CIFAR-10 and Tiny-ImageNet datasets.
4.5.3 Generative Model Type
We explored various generative model types, particularly focusing on path selection. To assess the impact of these paths on our method, we conducted experiments with three model types on the CIFAR-10 and Tiny-ImageNet datasets. Following [38], we quantified the straightness of various continuously differentiable trajectories. Interestingly, while all three model variants achieved comparable performance (Tab. 10), models with higher straightness required fewer pre-training steps to attain equivalent performance levels and demonstrated potential for superior outcomes.
Generative Model Type GVP ICFM OTCFM
CIFAR-10 (Accuracy) $97.35$ $97.59$ $97.65$
First to $97$% (Epoch) $162$ $135$ $128$
Straightness $7.36$ $0.34$ $0.11$
Tiny-ImageNet (Accuracy) $70.98$ $71.12$ $71.33$
First to $70$% (Epoch) $157$ $125$ $106$
Straightness $6.25$ $0.54$ $0.15$
Table 10: (Generative Model Type) Performance of different generative model types on the CIFAR-10 and Tiny-ImageNet datasets.
4.5.4 Scaling Up Network Parameters
We investigated the impact of scaling up U-Net models of varying network parameter sizes [48]. Table 11 provides a comparative analysis following fine-tuning on the CIFAR-10 and Tiny-ImageNet datasets. The findings suggest that increasing the model size results in a modest enhancement of final classification accuracy.
Model Param (M) CIFAR Acc. Tiny Acc.
PRG-ICFM-B $32$ $97.35$ $70.94$
PRG-ICFM-S $42$ $97.59$ $71.12$
PRG-ICFM-L $66$ $97.75$ $71.30$
PRG-ICFM-XL $122$ $97.88$ $71.80$
Table 11: (Scaling Capability) Performance gains of increasing U-Net model sizes on the CIFAR-10 and Tiny-ImageNet datasets.
5 Conclusion and Limitation
In this paper, we introduce a novel perspective on using pretrained reversible generative models as unsupervised visual representation learners. We systematically investigate the necessary designs of the two-stage pre-training and fine-tuning paradigm. Theoretical analysis supports the efficacy of the first-stage generative pre-training, while empirical verifications provide insights for designing the fine-tuning pipeline. Leveraging these simple yet effect techniques and findings, our method achieves state-of-the-art performance in image classification tasks using generative models. Additional experiments, such as those on OOD problems and transfer learning scenarios, further validate the robustness and other advantages of our approach. However, there remains room for improvement, including the integration of large foundation pretrained generative models from the open-source community (e.g., [52]) and optimizing the training epoch/speed for faster adaptation to downstream tasks.
Appendix A Why PRG is effective?
In the first pre-training stage, since no downstream task information is available, it remains unclear which features are most relevant. Consequently, no compression is performed during pre-training. Instead, the model aims to approximate the optimal representation as closely as possible in the absence of downstream information by leveraging unconditional flow matching to learn representations, effectively maximizing the lower bound of mutual information.
According to the manifold assumption [46, Chapter 20], data lies on a low-dimensional manifold $\mathcal{M}$ of intrinsic dimension $d^{*}$, which is significantly smaller than the ambient dimension $D$. Although, in theory, intermediate latents encode the same information as the original data, using these latents as inputs for downstream tasks is more meaningful. A well-trained flow model effectively extracts data from the low-dimensional manifold $\mathcal{M}$ and lifts it to the ambient space $\mathbb{R}^{D}$, ensuring that every sample in $\mathbb{R}^{D}$ remains semantically meaningful. In contrast, directly sampling from the original data space does not necessarily preserve semantic coherence. For instance, generating a $64\times 64$ image by sampling from a $64\times 64$ multivariate Gaussian distribution would typically result in a meaningless noise-like image resembling a snowflake pattern.
Moreover, since the flow trajectories of an ordinary differential equation (ODE) do not intersect [11], points within a given region remain confined, thereby preserving topological relationships throughout the transformation process.
Subsequent fine-tuning for downstream tasks is equally crucial. By adopting techniques such as optimal transport matching and gradient guidance, the model undergoes efficient adaptation, selectively discarding redundant information while selecting and enhancing task-critical features. As shown in Figure 4, mutual information decreases during fine-tuning, whereas task accuracy improves.
A.1 What’s the difference between fine-tuning and training a classifier based on the ODE architecture?
The effectiveness of fine-tuning depends significantly on whether the model undergoes pre-training beforehand. If a classifier is trained directly using the ODE architecture without pre-training, the model must learn meaningful intermediate latents solely through a supervised loss. This approach often fails, as the model struggles to discover good latent representations in the absence of prior knowledge. Our experiments (Sec. 4.5.1) also confirm that this method performs poorly.
In contrast, pre-training allows the model to extract a rich set of useful features as intermediate latents. During fine-tuning, the model can then selectively retain and enhance features relevant to downstream tasks while compressing redundant information in the latent space. This process leads to a more effective and structured representation, ultimately improving downstream task performance.
A.2 Model-agnosticity
Our method follows the model-agnostic principle as defined in Appendix D.5 of [56], where different architectures achieve comparable encoding quality through flow matching. This indicates that architectural constraints stem from implementation choices rather than the framework itself.
Notably, while early diffusion models primarily relied on U-Nets for practical stability, recent work [3] has demonstrated that flow matching can be successfully achieved with minimal architectural requirements, such as shallow skip connections. In this work, we achieve state-of-the-art performance across both U-Net (Sec. 4.4.1) and Transformer (Sec. 4.4.3) architectures, further validating the generality of our approach.
A.3 Infinite-Layer Expressiveness
The infinite-layer extractor boosts accuracy by extending training rather than inference steps (Fig.6: CIFAR 0.46 vs. 0.97, Tiny 0.19 vs. 0.71). However, gains plateau after a certain point, with more complex datasets requiring longer training for optimal performance. Meanwhile, the optimal flow-matching setup achieves full experimental coverage in just $5$ inference steps. Full evaluation takes 3 seconds for CIFAR and 9 s for the Tiny-ImageNet testing set.